Pollution is expensive to clean up. The cost to clean up p% of pollution is modeled by:
C=12000/400-4p How much does it cost to remove 80% of pollution

Answer by fractalier(6550)

I think you mean
C=12000/(400-4p)
Then, if so, merely plug in 80 for p and get
C=12000/(400-4*80)
C=12000/(400-320)
C=12000/80
C=150